Assess the three different models used to value stocks based on different dividend patterns. List the different models and provide an example of each.
Answer #1

The three different models used to value stocks based on different dividend patterns are

1. Zero-growth Dividend Model- In this dividend have a growth rate of zero which assumes that all dividends paid by a stock remain the same. There is no growth in dividends. Under this model the stock price is equal to the annual dividends divided by the required rate of return.

The formula used for estimating the stock value is

Intrinsic value of the stock = Annual Dividends/ Required Rate of Return which is the cost of equity

i.e I=D/K

eg. If a preferred share of stock pays dividends of $2 per year, and the required rate of return for the stock is 7%, then what is its intrinsic value

Using the zero growth dividend discount model formula, Intrinsic Value = $2/0.07 = $28.57

2. Constant-Growth Dividend Model -In this dividends grow at a constant rate, which assumes that dividends grow by a specific percent annually.

The formula used for estimating the stock value is

Intrinsic value of the stock = Current dividend (1+ Dividend Growth Rate)/ (Required Rate of Return which is the cost of equity - Dividend Growth Rate)

i.e I = D0 (1+G)/(R-G)

I = D1/(R-G)

Under Constant-Growth Dividend Model, dividends for some firms grow at a steady rate.

e.g. If a stock pays a $3 dividend this year, and the dividend has been growing 5% annually, then what will be the intrinsic value of the stock , assuming a required rate of return of 10%

Using Constant-Growth Dividend Model,

Stock Value= ($3 x 1.05)/(0.10-0.05)=3.15/0.05

=$63

3. Mixed Growth Dividend Model- In this dividends have a mixed growth rate pattern. Under this model, growth is divided into two or three phases. It is also known as Variable Growth Dividend Discount Model. It has a different rates of growth, firstly an initial high rate of growth, then a transition to slower growth, and lastly, a steady rate of growth.